# Tricriticality and finite-size scaling in the triangular Blume-Capel ferromagnet

### v1

@@ III.3.2 Second-order transition regime
with $\{b_{\Delta_{O}},b^{\prime}_{\Delta_{O}}\}$ non-universal constants (similarly for the fitting coefficients $\{b^{\prime}_{m,(n)},b_{C},b^{\prime}_{C},b^{\prime}_{\chi}\}$ used in the scaling Eqs. (32) - (34) below) and $\omega$ the universal corrections-to-scaling exponent, fixed to the Ising value $7/4$ (see, for example, the discussion in the supplementary material of Ref. [77] and comment [78]). Simultaneously fitting the numerical data for the larger system sizes studied ($L_{\rm min}=32$) to the power-law (31) we obtain $\Delta_{\rm c}=1.014(1)$ (see also Table 1) and $\nu=1.00(2)$, the latter in very good agreement with the exact value $\nu=1$ of the Ising ferromagnet. Figure 14 yields a supplementary estimation of the correlation-length’s exponent $\nu$ via the finite-size scaling behavior of the logarithmic derivatives of the order parameter. In particular, we demonstrate in Fig. 14 the cases with $n=2$ and $3$ of Eq. (26). Fittings of the form [66]
$C^{\ast}_{L}=b_{C}+b^{\prime}_{C}\ln{(L)},$ (33)
$\chi^{\ast}_{L}=L^{\gamma/\nu}(1+b^{\prime}_{\chi}L^{-\omega}).$ (34)
@@ Acknowledgements.
We would like to dedicate this paper to the memory of Walter Selke who passed away in December 2023 and whose seminal contributions shaped the field of Statistical Physics. N.G.F. would like to thank Nigel Wilding for some very useful comments and suggestions at the final stage of this work. Part of the numerical calculations reported in this paper were performed at the High-Performance Computing cluster CERES of the University of Essex. The work of A.V. and N.G.F. was supported by the Engineering and Physical Sciences Research Council (grant EP/X026116/1 is acknowledged). D.H.K. acknowledges the support from GIST Research Institute (GRI) grant funded by the GIST.

### v2

@@ III.3.2 Second-order transition regime
with $\{b_{\Delta_{O}},b^{\prime}_{\Delta_{O}}\}$ non-universal constants (similarly for the fitting coefficients $\{b^{\prime}_{m,(n)},b_{C},b^{\prime}_{C},b^{\prime}_{\chi}\}$ used in the scaling Eqs. (32) - (34) below) and $\omega$ the universal corrections-to-scaling exponent, fixed to the Ising value $7/4$ (see, for example, the discussion in the supplementary material of Ref. [83] and comment [84]). Simultaneously fitting the numerical data for the larger system sizes studied ($L_{\rm min}=32$) to the power-law (31) we obtain $\Delta_{\rm c}=1.014(1)$ (see also Table 1) and $\nu=1.00(2)$, the latter in very good agreement with the exact value $\nu=1$ of the Ising ferromagnet. Figure 14 yields a supplementary estimation of the correlation-length’s exponent $\nu$ via the finite-size scaling behavior of the logarithmic derivatives of the order parameter. In particular, we demonstrate in Fig. 14 the cases with $n=1$ and $2$ of Eq. (26). Fittings of the form [69]
$C^{\ast}_{L}=b_{C}+b^{\prime}_{C}\ln{(L)},$ (33)
$\chi^{\ast}_{L}=L^{\gamma/\nu}(1+b^{\prime}_{\chi}L^{-\omega}).$ (34)
@@ Acknowledgements.
N.G.F. would like to thank Walter Selke, who sadly passed away in December 2023, for his ongoing collaboration on the Blume-Capel model. We are grateful to Nigel Wilding for some very useful comments and suggestions at the final stage of this work. Part of the numerical calculations reported in this paper were performed at the High-Performance Computing cluster CERES of the University of Essex. The work of A.V. and N.G.F. was supported by the Engineering and Physical Sciences Research Council (grant EP/X026116/1 is acknowledged). D.H.K. acknowledges the support from GIST Research Institute (GRI) grant funded by the GIST and appreciates APCTP and KIAS for their hospitality during the completion of this work.